​How many moles of LiCl are required to make 1200 g of a 1.3% (w:w) solution?
Answer #1

Concentration LiCl = 1.3 % (w/w)

Concentration LiCl = 1.3 / 100 (w/w)

Concentration LiCl = 0.013 (w/w)

mass of solution = 1200 g

mass LiCl = (Concentration LiCl) * (mass of solution)

mass LiCl = (0.013) * (1200 g)

mass LiCl = 15.6 g

moles LiCl = (mass LiCl) / (molar mass LiCl)

moles LiCl = (15.6 g) / (42.394 g/mol)

moles LiCl = 0.37 mol